# Superficial Self-Improved Reasoners Benefit from Model Merging

### v1

@@ Abstract
As scaled language models (LMs) approach human-level reasoning capabilities, self-improvement emerges as a solution to synthesizing high-quality data corpus. While previous research has identified model collapse as a risk in self-improvement, where model outputs become increasingly deterministic, we discover a more fundamental challenge: the superficial self-improved reasoners phenomenon. In particular, our analysis reveals that even when LMs show improved in-domain (ID) reasoning accuracy, they actually compromise their generalized reasoning capabilities on out-of-domain (OOD) tasks due to memorization rather than genuine learning. Through a systematic investigation of LM architecture, we discover that during self-improvement, LM weight updates are concentrated in less reasoning-critical layers, leading to superficial learning. To address this, we propose Iterative Model Merging (IMM), a method that strategically combines weights from original and self-improved models to preserve generalization while incorporating genuine reasoning improvements. Our approach effectively mitigates both LM collapse and superficial learning, moving towards more stable self-improving systems.
@@
Xiangchi Yuan1, Chunhui Zhang2, Zheyuan Liu3, Dachuan Shi1, Soroush Vosoughi2, Wenke Lee1 1Georgia Institute of Technology, 2Dartmouth College, 3University of Notre Dame {xyuan300, dshi77, wenke}@gatech.edu {chunhui.zhang.gr, soroush}@dartmouth.edu zliu29@nd.edu
@@ 1 Introduction
The remarkable reasoning capabilities Jaech et al. (2024); Zhang et al. (2024); Guo et al. (2025) of large language models (LLMs) largely benefits from vast amounts of high-quality reasoning data. However, as the data corpus runs out Sutskever (2024) and increasingly powerful models approach human-level intelligence DeepMind (2024a, b), pressing issues emerge: (i) How to advance models’ reasoning capabilities despite data scarcity? (ii) How to obtain training data that exceeds human-level performance for next-generation models? A promising answer to both questions is model self-improvement or self-evolution, where models autonomously generate infinite high-quality data, which potentially surpasses human annotations, to continuously enhance their own performance.
@@
Although self-improvement has achieved remarkable success in specific domains such as mathematics OpenAI (2025); DeepMind (2024a), coding Li et al. (2022), and games Hu et al. (2024); Silver et al. (2018), recent studies reveal significant risks associated with using self-generated synthetic data for fine-tuning: in particular, model performance can degrade over multiple iterations of self-improvement, a phenomenon known as model collapse. Shumailov et al. (2023). In current research, model collapse is primarily attributed to a reduction in sampling diversity Shumailov et al. (2023); Alemohammad et al. (2024); Guo et al. (2024). To mitigate this problem, several studies suggest refreshing synthetic data with real data Bertrand et al. (2024); Alemohammad et al. (2024), accumulating data across training steps Gerstgrasser et al. (2024), and incorporating data verifiers Gillman et al. (2024) or correctors Feng et al. (2025). However, by focusing solely on data quality and diversity, these approaches overlook a more critical question: whether self-improvement genuinely enhances reasoning capabilities or merely memorizes the training distribution. This distinction becomes crucial when considering the model’s ability to generalize beyond its training data.
In this paper, we investigate a fundamental risk in model self-improvement for reasoning tasks that deepens the known challenge of model collapse. We identify a phenomenon we call Superficial Self-Improved Reasoners, where models appear to improve but actually fail to develop genuine reasoning capabilities. While these models show enhanced performance on in-domain (ID) reasoning tasks, they significantly underperform on out-of-domain (OOD) tasks, suggesting memorization rather than genuine reasoning improvement. To understand the mechanistic cause of this phenomenon, we perform a systematic analysis of the model architecture during self-improvement. By examining layer importance and parameter changes, we uncover a critical mismatch: the largest weight updates occur in layers that contribute least to reasoning, while reasoning-critical layers receive minimal updates. This mismatch explains why models tend to memorize training patterns rather than develop generalizable reasoning skills. To address this issue, we propose Iterative Model Merging (IMM), a novel method that strategically combines weights from original and self-improved models. IMM specifically targets the layer misalignment problem by preserving the stability of reasoning-critical layers while allowing beneficial updates from self-improvement. As demonstrated in Figure 1, this approach effectively balances performance improvements with preserved generalization ability.
@@
  • We provide an explanation for the emergence of Superficial Self-Improved Reasoners by highlighting a mismatch between the reasoning-critical layers and the layers that undergo the largest weight changes during self-improvement.
  • We propose IMM (Iterative Model Merging) to mitigate Superficial Self-Improved Reasoners. IMM offers an efficient approach to integrate the reasoning improvements of the self-improved model while preserving the generalization ability of the original model.
@@ 2 Related Work
LLMs for Reasoning
LLMs have demonstrated remarkable success across various reasoning tasks, including mathematical problem-solving, code generation, common-sense reasoning, and multimodal understanding Yu et al. (2024b); Lewkowycz et al. (2022); Wang et al. (2023); Zhang et al. (2025a, b). Beyond leveraging sophisticated prompting techniques to enhance reasoning capabilities Kojima et al. (2022); Wei et al. (2022); Zheng et al. (2023); Yao et al. (2024), many methods focus on fine-tuning LLMs with reasoning datasets to create more robust reasoners Lu et al. (2024); Yu et al. (2024b). For instance, approaches like SI Huang et al. (2022), STaR Zelikman et al. (2022), V-STaR Hosseini et al. (2024), and rSTaR Qi et al. (2025) fine-tune LLMs on task-specific datasets or synthesize reasoning data tailored for corresponding tasks. In addition to training models to generate correct answers, some studies introduce external verifiers Cobbe et al. (2021); Lightman et al. (2023); Hosseini et al. (2024); Yuan et al. (2024a) that select the best answer from a set of candidate solutions.
@@
As real-world data becomes increasingly scarce Sutskever (2024), synthetic data is playing a crucial role in training modern generative models due to its low cost and infinite availability. However, recent studies have revealed the risks associated with this "free lunch," a phenomenon known as model collapse Shumailov et al. (2023). The model collapse has been extensively identified and analyzed in both computer vision Hataya et al. (2023); He et al. (2022); Bohacek and Farid (2023) and natural language processing Alemohammad et al. (2024); Liu et al. (2024); Gerstgrasser et al. (2024). Researchers have investigated its underlying causes from both empirical Padmakumar and He (2024); Dou et al. (2024); Guo et al. (2023) and theoretical perspectives Yuan et al. (2024a); Bertrand et al. (2023); Seddik et al. (2024). Current approaches to mitigating model collapse predominantly focus on data-centric methods. Feng et al. (2025) show that imperfect verifiers can help prevent model collapse by selecting appropriate data. Shumailov et al. (2023) proposes mixing data from previous iterations to prevent performance degradation, while Gerstgrasser et al. (2024) demonstrates that accumulating synthetic data over iterations reduces the risk of collapse.
@@ 3.1 Identify Superficial Self-improved Reasoners from OOD datasets
In this part, we identify Superficial Self-improved Reasoners by self-improving LLMs on the in-domain (ID) reasoning datasets and test them on out-of-domain (OOD) reasoning datasets.
@@
We begin by establishing the self-improvement framework through the generation of reasoning data. Following prior work Zelikman et al. (2022), we first synthesize reasoning data for fine-tuning. Let $\mathcal{D}=\{(q_{i},a_{i})\}_{i=1}^{n_{d}}$ denote a training dataset containing $n_{d}$ reasoning questions $q_{i}$ and corresponding final answers $a_{i}$. We also use Chain-of-Thought prompting Wei et al. (2022) with in-context examples in this process (details in Appendix A). In the second step, we sample multiple solutions for each $q_{i}$ using non-zero sampling temperatures, resulting in a synthetic dataset $\mathcal{D}_{S}=\{(q_{i},\{(\hat{r}_{ij},\hat{a}_{ij})\}_{j=1}^{k})\}$, where $k$ represents the number of sampled solutions. Here, $\hat{r}_{ij}$ denotes the $j$-th reasoning path (i.e., rationale) generated by the model for $q_{i}$, and $\hat{a}_{ij}$ is the model’s corresponding final answer. Incorrect solutions are then filtered out by comparing the sampled answers $\hat{a}_{ij}$ with the ground-truth answers $a_{i}$. Finally, we fine-tune the model on the filtered dataset $\tilde{\mathcal{D}}_{G}$ using supervised fine-tuning (SFT) to maximize the likelihood of generating reasoning paths $r$, optimizing the following objective:
@@
After applying the self-improvement framework to LLMs of various scales on in-domain (ID) datasets, we evaluate their performance on out-of-domain (OOD) reasoning datasets. The results, presented in Figure 2, reveal that while self-improvement enhances reasoning performance on ID datasets, it leads to a noticeable decline in performance on OOD datasets. This phenomenon suggests that although self-improvement improves metrics on ID reasoning tasks, it fails to enhance generalized reasoning capabilities and may even degrade them. We refer to this behavior as the emergence of Superficial Self-Improved Reasoners.
@@ 3.2 Investigating the Causes of Superficial Self-Improved Reasoners
We leverage this layer importance score $I^{n}$ to identify which layers contribute most significantly to reasoning tasks. As illustrated in Figure 3, the middle layers are less important while the early and late layers are more important for the reasoning (MATH) task. However, for the non-reasoning dataset BookCorpus, the middle layers are more important. This observation highlights the early and late layers as reasoning-critical layers, distinguishing their specialized function in reasoning.
@@
After fine-tuning the LLMs on reasoning data, the weights are updated, enabling the model to learn reasoning capabilities. We now analyze these weight changes. Let $\Delta\mathbf{W}^{n}$ represent the total weight change at the $n$-th layer after supervised fine-tuning (SFT), defined as:
@@
where $\mathbf{W}^{k,n}$ denotes the original $k$-th weight matrix and $\mathbf{W}^{k,n}_{\text{SFT}}$ is the fine-tuned weight matrix. Figure 4 illustrates the weight change $\Delta\mathbf{W}^{n}$ across different layers. For the self-improved model, the largest weight change occurs in the middle layers. In contrast, for the math model which is fully post-trained with stronger generalized reasoning capability, the most significant weight changes are concentrated in the early and late layers.
@@
This observation suggests that during self-improvement, the model does not exclusively enhance its reasoning ability but also exhibits a tendency to overfit the training data, effectively "memorizing" it. This overfitting behavior explains the improved performance on ID datasets while compromising the model’s generalization to OOD tasks. The performance comparison in Figure 2 further supports this conclusion. We summarize all experimental findings in Table 1, which leads to the following key insights: (i) during self-improvement on reasoning tasks, LLMs may show improved reasoning performance on ID tasks but lose generalized reasoning ability on OOD tasks; (ii) This phenomenon arises from a mismatch between the reasoning-critical layers and the layers with significant weight changes, suggesting that the model memorizes the training data rather than truly learning generalized reasoning capability.
@@ 4 Superficial Self-improved Reasoners Benefit from Iterative Model Merging
Figure 5: The overall framework. (a) The model generates chain-of-thought (CoT) answers for the given questions, and incorrect answers are filtered out using the ground-truth. The remaining correct answers are used for supervised fine-tuning (SFT) to self-improve the model. (b) IMM iteratively SFT the model and merges the self-improved models with the base model to balance reasoning enhancement and generalization.
@@
$\boldsymbol{\delta^{t}}=\begin{cases}\boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta_{m}^{t}},&\text{if }t>0\text{, SFT merged model},\\ \boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta},&\text{if }t=0\text{, SFT base model}.\end{cases}$ (5)
@@
$\boldsymbol{\theta_{m}^{t+1}}=\alpha\boldsymbol{\theta}+(1-\alpha)(\boldsymbol{\theta^{t}}+\boldsymbol{\hat{\delta^{t}}}),$ (6)
where $\alpha$ is a scaling parameter that controls the balance between the base model weights and the self-improved model weights.
@@ 5 Experiments
Figure 6: The model performances on in-domain (ID) datasets performance. SFT n and Merge n denote the SFT model and merged model in the n-th iteration cycle. The model collapse happens from the first or second iteration for baselines while our method avoids it and achieves the best performance after model merging.
@@
Figure 8: ID performance with different k for scaling up test-time-computing Pass@k.

### v2

@@ Abstract
As scaled language models (LMs) approach human-level reasoning capabilities, self-improvement emerges as a solution to synthesizing high-quality data corpus. While previous research has identified model collapse as a risk in self-improvement, where model outputs become increasingly deterministic, we discover a more fundamental challenge: the superficial self-improved reasoners phenomenon. In particular, our analysis reveals that even when LMs show improved in-domain (ID) reasoning accuracy, they actually compromise their generalized reasoning capabilities on out-of-domain (OOD) tasks due to memorization rather than genuine learning. Through a systematic investigation of LM architecture, we discover that during self-improvement, LM weight updates are concentrated in less reasoning-critical layers, leading to superficial learning. To address this, we propose Iterative Model Merging (IMM), a method that strategically combines weights from original and self-improved models to preserve generalization while incorporating genuine reasoning improvements. Our approach effectively mitigates both LM collapse and superficial learning, moving towards more stable self-improving systems. Code is available [^1] Code is available at IMM..
@@
Xiangchi Yuan1, Chunhui Zhang2, Zheyuan Liu3, Dachuan Shi1, Leyan Pan1, Soroush Vosoughi2, Wenke Lee1 1Georgia Institute of Technology, 2Dartmouth College, 3University of Notre Dame {xyuan300, dshi77, leyanpan}@gatech.edu, wenke@cc.gatech.edu {chunhui.zhang.gr, soroush}@dartmouth.edu zliu29@nd.edu
@@ 1 Introduction
The reasoning capabilities Jaech et al. (2024); Zhang et al. (2024); Guo et al. (2025) of large language models (LLMs) largely benefits from vast amounts of high-quality reasoning data. However, as the data corpus runs out Sutskever (2024) and increasingly powerful models approach human-level intelligence DeepMind (2024a, b), pressing issues emerge: (i) How to advance models’ reasoning capabilities despite data scarcity? (ii) How to obtain training data that exceeds human-level performance for next-generation models? A promising answer to both questions is model self-improvement or self-evolution, where models autonomously generate infinite high-quality data, which potentially surpasses human annotations, to continuously enhance their own performance.
@@
Although self-improvement has achieved remarkable success in specific domains such as mathematics OpenAI (2025); DeepMind (2024a), coding Li et al. (2022), and games Hu et al. (2024); Silver et al. (2018), recent studies reveal significant risks associated with using self-generated synthetic data for fine-tuning: in particular, model performance can degrade over multiple iterations of self-improvement, a phenomenon known as model collapse. Shumailov et al. (2023). In current research, model collapse is primarily attributed to a reduction in sampling diversity Shumailov et al. (2023); Alemohammad et al. (2024); Guo et al. (2024b). To mitigate this problem, several studies suggest refreshing synthetic data with real data Bertrand et al. (2024); Alemohammad et al. (2024), accumulating data across training steps Gerstgrasser et al. (2024), and incorporating data verifiers Gillman et al. (2024) or correctors Feng et al. (2025). However, by focusing solely on data quality and diversity, these approaches overlook a more critical question: whether self-improvement genuinely enhances reasoning capabilities or merely memorizes the training distribution. This distinction becomes crucial when considering the model’s ability to generalize beyond its training data.
In this paper, we investigate a risk in model self-improvement for reasoning tasks that deepens the known challenge of model collapse. We identify a phenomenon we call Superficial Self-Improved Reasoners, where models appear to improve but actually fail to develop genuine reasoning capabilities. While these models show enhanced performance on in-domain (ID) reasoning tasks, they significantly underperform on out-of-domain (OOD) tasks, suggesting memorization rather than genuine reasoning improvement. To understand the mechanistic cause of this phenomenon, we perform a systematic analysis of the model architecture during self-improvement. By examining layer importance and parameter changes, we uncover a critical mismatch: the largest weight updates occur in layers that contribute least to reasoning, while reasoning-critical layers receive minimal updates. This mismatch explains why models tend to memorize training patterns rather than develop generalizable reasoning skills. To address this issue, we propose Iterative Model Merging (IMM), a novel method that strategically combines weights from original and self-improved models. IMM specifically targets the layer misalignment problem by preserving the stability of reasoning-critical layers while allowing beneficial updates from self-improvement. As demonstrated in Figure 1, this approach effectively balances performance improvements with preserved generalized reasoning capability.
@@
  • We provide an explanation for this phenomenon by highlighting a mismatch between the reasoning-critical layers and the layers that undergo the largest weight changes.
  • We propose IMM to mitigate this phenomenon. IMM offers a simple, general, and effective approach to integrate the reasoning improvements of the self-improved model while preserving the generalization of the original model.
@@ 2 Related Work
As real-world data becomes increasingly scarce Sutskever (2024), synthetic data is playing a crucial role in training modern generative models due to its low cost and infinite availability. However, recent studies have revealed the risks associated with this "free lunch," a phenomenon known as model collapse Shumailov et al. (2023). The model collapse has been extensively identified and analyzed in both computer vision Hataya et al. (2023); He et al. (2022); Bohacek and Farid (2023) and natural language processing Alemohammad et al. (2024); Gerstgrasser et al. (2024). Researchers have investigated its underlying causes from both empirical Padmakumar and He (2024); Guo et al. (2023) and theoretical perspectives Yuan et al. (2024); Bertrand et al. (2023); Seddik et al. (2024); Fu et al. (2024a). Current approaches to mitigating model collapse predominantly focus on data-centric methods. Feng et al. (2025) show that imperfect verifiers can help prevent model collapse by selecting appropriate data. Shumailov et al. (2023) proposes mixing data from previous iterations to prevent performance degradation, while Gerstgrasser et al. (2024) demonstrates that accumulating synthetic data over iterations reduces the risk of collapse.
Appendix C.9 discusses additional related works on LLM for reasoning. The connection with catastrophic forgetting is discussed in Appendix C.3.
@@ 3.1 Identify Superficial Self-improved Reasoners from OOD datasets
In this part, we identify Superficial Self-improved Reasoners by self-improving LLMs on the ID reasoning datasets and test them on OOD datasets.
@@
We begin by establishing the self-improvement framework through the generation of reasoning data. Following prior work Zelikman et al. (2022), we first synthesize reasoning data for fine-tuning. Let $\mathcal{D}=\{(q_{i},a_{i})\}_{i=1}^{n_{d}}$ denote a training dataset containing $n_{d}$ reasoning questions $q_{i}$ and corresponding final answers $a_{i}$. We also use Chain-of-Thought prompting Wei et al. (2022) in this process (details in Appendix A.1). In the second step, we sample multiple solutions for each $q_{i}$ using non-zero sampling temperatures, resulting in a synthetic dataset $\mathcal{D}_{S}=\{(q_{i},\{(\hat{r}_{ij},\hat{a}_{ij})\}_{j=1}^{k})\}$, where $k$ represents the number of sampled solutions. Here, $\hat{r}_{ij}$ denotes the $j$-th reasoning path (i.e., rationale) generated by the model for $q_{i}$, and $\hat{a}_{ij}$ is the model’s corresponding final answer. Incorrect solutions are then filtered out by comparing the sampled answers $\hat{a}_{ij}$ with the ground-truth answers $a_{i}$. Finally, we fine-tune the model on the filtered dataset $\tilde{\mathcal{D}}_{G}$ using supervised fine-tuning (SFT) to maximize the likelihood of generating reasoning paths $r$, optimizing the following objective:
@@
After applying the self-improvement framework to LLMs of various scales on ID datasets, we evaluate their performance on OOD reasoning datasets. The results, presented in Figure 2, reveal that while self-improvement enhances reasoning performance on ID datasets, it leads to a noticeable decline in performance on OOD datasets. This phenomenon suggests that although self-improvement improves metrics on ID reasoning tasks, it fails to enhance generalized reasoning capabilities and may even degrade them. We refer to this behavior as the emergence of Superficial Self-Improved Reasoners.
@@ 3.2 Investigating the Causes of Superficial Self-Improved Reasoners
We leverage this layer importance score $I^{n}$ to identify which layers contribute most significantly to reasoning tasks. As illustrated in Figure 3, the middle layers are less important while the early and late layers are more important for the reasoning (MATH) tasks. We also find similar performance on code reasoning tasks, as illustrated in Appendix 11. However, for the non-reasoning dataset BookCorpus, the middle layers are more important. This observation highlights the early and late layers as reasoning-critical layers (More clarification for this term is in Appendix C.4), distinguishing their specialized function in reasoning.
@@
After fine-tuning the LLMs on reasoning data, the weights are updated, enabling the model to learn reasoning capabilities. We now analyze these weight changes. Let $\Delta\mathbf{W}^{n}$ represent the total weight change at the $n$-th layer after SFT:
@@
where $\mathbf{W}^{k,n}$ denotes the original $k$-th weight matrix and $\mathbf{W}^{k,n}_{\text{SFT}}$ is the fine-tuned weight matrix. Figure 4 illustrates the weight change $\Delta\mathbf{W}^{n}$ across different layers. For the self-improved model, the largest weight change occurs in the middle layers. In contrast, for the math model which is fully post-trained with stronger generalized reasoning capability, the most significant weight changes are concentrated in the early and late layers. A similar condition happens for real data with limited training data size, as analyzed in Appendix B.1.
Figure 5: The overall framework: (a) The model generates chain-of-thought (CoT) answers for the given questions, and incorrect answers are filtered out using the ground-truth. The remaining correct answers are used for SFT to self-improve the model. (b) IMM iteratively SFT the model and merges the self-improved models with the base model to balance reasoning enhancement and generalization.
@@
This observation suggests that during self-improvement, the model does not exclusively enhance its reasoning ability but also exhibits a tendency to overfit the training data, effectively "memorizing" it. This overfitting behavior explains the improved performance on ID datasets while compromising the model’s generalization to OOD tasks. The performance comparison in Figure 2 further supports this conclusion. We summarize all experimental findings in Table 1, which leads to the following key insights: (i) during self-improvement on reasoning tasks, LLMs may show improved reasoning performance on ID tasks but lose generalized reasoning ability on OOD tasks; (ii) This phenomenon arises from a mismatch between the reasoning-critical layers and the layers with significant weight changes, suggesting that the model memorizes the training data rather than truly learning generalized reasoning capability. We further provide analysis on the reasons for this mismatch phenomenon in Appendix C.2.
@@ 4 Superficial Self-improved Reasoners Benefit from Iterative Model Merging
$\boldsymbol{\delta^{t}}=\begin{cases}\boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta_{m}^{t-1}},&\text{if }t>0\text{, SFT }\boldsymbol{\theta_{m}^{t-1}},\\ \boldsymbol{\theta^{t}_{SFT}}-\boldsymbol{\theta},&\text{if }t=0\text{, SFT }\boldsymbol{\theta}.\end{cases}$ (5)
@@
[…]
where $\alpha$ is a scaling parameter that controls the balance between the base model weights and the self-improved model weights. Although we use a uniform $\alpha$ for all layers, which makes reasoning-critical layers’s weight change remain minimal at the first iteration, this generalized way makes the model avoid overfitting and learn the generalized reasoning capability, which makes reasoning-critical layers’ weights change increase more compared to the reasoning-trivial layers in the next iterations to learn generalized reasoning capability, as analyzed in Appendix B.9. The overall merging strategy is scalable for multiple iterations and larger models, with complexity analysis presented in Appendix B.10.
@@
Importance-based Iterative Model Merging (IIMM)
We also propose IIMM, which is motivated to aggressively merge the model according to the layer importance as follows:
[…]
where $n$ denotes $n$-th layer of the model with $N$ layers. However, we find that IIMM is outperformed by IMM because of instability and overfitting datasets for importance score calculation. The detailed experiment and analysis are provided in Appendix B.5.
@@ 5 Experiments
Figure 6: The model performances on in-domain (ID) datasets. SFT n and Merge n denote the SFT model and merged model in the n-th iteration cycle. The model collapse happens from the first or second iteration for baselines, while our method avoids it and achieves the best performance after model merging.
@@
Figure 8: ID performance with different k for scaling up test-time-computing Pass@k on GSM8K.